OpenStudy (anonymous):

an orchard has 2030 trees. the number of trees in each row exceeds twice the number of rows by 12. how many trees are in each row?

OpenStudy (anonymous):

Let t and r be (the number of trees per row) and (the number of rows) respectively.\[\{r* t=2030,t=2* r+12\} \]\[\{t=70,r=29\} \]

OpenStudy (anonymous):

can you please show your work, im still clueless to how you got 70 and 29